OpenStudy (anonymous):

answer check? Write an equation for the line that contains the point (-6, -8) and is parallel to the line 3x – 2y = -12. A. -3x + 2y = -10 B. -3x + 2y = 2 C. -3x + 2y = 22 D. 2x + 3y = -36 i got A

OpenStudy (danjs):

B

OpenStudy (danjs):

Plug in (-6,-8) into A and see if it is -10

jimthompson5910 (jim_thompson5910):

Let's check choice A -3x + 2y = -10 -3(-6) + 2(-8) = -10 ... plug in (x,y) = (-6,-8) 18 - 16 = -10 2 = -10 ... this is false so (-6,-8) is not on this line

OpenStudy (anonymous):

ohh i got it now thanks jim

jimthompson5910 (jim_thompson5910):

np